# A Spatio-temporal CP decomposition analysis of New England region in the US

### v2

@@ 4.1 STPCA-CPD using Alternating Least Square method
In the following experiments we applied the STPCA-CPD algorithm to the real dataset describe in section˜2. We did the decomposition varying the number of components (i.e. the rank). We also compare the performance of our algorithm with the well known HOSVD+CPD where the initialization of the factor matrices is done using Higher Order Singular Value Decomposition (HOSVD), and a random initialization coupled with using ALS to find the CP factor matrices. We compute the reconstruction error using the relative error for each method. All the decompositions for HOSVD+CPD, STPCA+CPD, and random+CPD were computed using the matlab tensor toolbox (Kolda and Bader, 2009; Bader et al., 2024) and the tensorLab toolbox De Lathauwer et al. (2023).
@@ 4.2 Clustering analysis of the CP factors per mode
To continue this analysis we are now looking at the K-means clustering (see section˜3.3) of the CPD factors with the initialization methods used above. The goal in this case is to use the components from all the method in mode 1 (temporal mode) and mode 2 (spatial mode), and do a K-means clustering analysis. We use the silhouette score as described in section˜3.3. The k-means analysis run through a range of k from 2 to 12 for all the decompositions. Table 2 shows the results for rank 2 and Table 3 shows the results for rank 3. In both cases, STPCA+CPD outperform the other methods.
@@ 5 Conclusions
In this study, we employed a Spatio-Temporal Principal Component Analysis (STPCA) to initialize the factors in our CANDECOMP/PARAFAC (CP) decomposition, which was subsequently optimized using the Alternating Least Squares (ALS) algorithm. The numerical experiments yielded promising results, demonstrating a distinct advantage in leveraging the spatio-temporal dependencies inherent in the data into the CP factorization process. Model performance was assessed through the computation of the relative reconstruction error for each configuration.
Furthermore, we applied k-means clustering to the extracted factor matrices and evaluated clustering quality using silhouette scores. The proposed STPCA+CPD approach consistently outperformed baseline methods, confirming the benefits of our initialization strategy.
For future research, this framework could be extended by integrating deep learning architectures to enhance its capacity to capture complex spatio-temporal dynamics. Additionally, we plan to investigate more robust spatio-temporal tensor decomposition schemes and to generalize the tensor framework for forecasting and downscaling applications.
6 Acknowledgments
The author would like to thank Professor Julie Carreau for helpful discussion concerning this project.
Code availability section
All data is publicly available and cited. The function stpca_to_cp_init in cp_init.R is for the initialization and is available on github.
Contact: fsanogo@bates.edu, (207) 786-6144.
Hardware requirements: No specialized hardware is required. All analyses were performed on standard personal computing hardware with sufficient RAM to handle the data size and MATLAB/R software installed.
Program language: The R and Matlab programming languages were used.
Software required: The analysis primarily utilizes the Tensor Toolbox for MATLAB (Kolda and Bader) and TensorLab (De Lathauwer et al.) within the MATLAB environment. Specific custom for the initialization was written in R. The source codes are available for downloading at the link:
https://github.com/fatsanogo/STPCA-CP-Decomposition-ALS

### v3

@@ 4.1 STPCA-CPD using Alternating Least Square method
In the following experiments we applied the STPCA-CPD algorithm to the real dataset describe in section 2. We did the decomposition varying the number of components (i.e. the rank). We also compare the performance of our algorithm with the well known HOSVD+CPD where the initialization of the factor matrices is done using Higher Order Singular Value Decomposition (HOSVD), and a random initialization coupled with ALS to find the CP factor matrices. We compute the reconstruction error using the relative error for each method. All the decompositions for HOSVD+CPD, STPCA+CPD, and random+CPD were computed using the matlab tensor toolbox [doi:10.1137/07070111X, tensor_toolbox_2024] and the tensorLab toolbox [TensorLab].
@@ 4.2 Clustering analysis of the CP factors per mode
To continue this analysis we are now looking at the K-means clustering (see section 3.3) of the CPD factors with the initialization methods used above. The goal in this case is to use the components from all the method in mode 1 (temporal mode) and mode 2 (spatial mode), and do a K-means clustering analysis. We use the silhouette score as described in section 3.3. The k-means analysis run through a range of k from 2 to 12 for all the decompositions. Table 2 shows the results for rank 2 and Table 3 shows the results for rank 3. In both cases, STPCA+CPD outperform the other methods.
@@ 5 Conclusions
In this study, we employed a Spatio-Temporal Principal Component Analysis (STPCA) to initialize the factors in our CANDECOMP/PARAFAC (CP) decomposition, which was subsequently optimized using the Alternating Least Squares (ALS) algorithm. The numerical experiments yielded promising results, demonstrating a distinct advantage in leveraging the spatio-temporal dependencies inherent in the data into the CP factorization process. Model performance was assessed through the computation of the relative reconstruction error for each configuration. Furthermore, we applied k-means clustering to the extracted factor matrices and evaluated clustering quality using silhouette scores. The proposed STPCA+CPD approach consistently outperformed baseline methods, confirming the benefits of our initialization strategy. For future research, this framework could be extended by integrating deep learning architectures to enhance its capacity to capture complex spatio-temporal dynamics. Additionally, we plan to investigate more robust spatio-temporal tensor decomposition schemes and to generalize the tensor framework for forecasting and downscaling applications.
[…]
Supplementary information
[…]
AcknowledgementsThe author would like to thank Professor Julie Carreau for helpful discussion concerning this project.
Declarations
  • The author declares that no funds, grants, or other support were received during the preparation of this manuscript.
  • The author hereby declare that they have no known competing financial interests or personal relationships that could have potentially influenced the work presented in this paper.
  • All data is publicly available and cited.
  • No materials was used
  • Codes available are available upon request.
  • All work in this manuscript was done by the author.